Charles Lackey operates a bakery in Idaho​ Falls, Idaho. Because of its excellent product and excellent​ location, demand has increased by 35​% in the last year. On far too many​ occasions, customers have not been able to purchase the bread of their choice. Because of the size of the​ store, no new ovens can be added. At a staff​ meeting, one employee suggested ways to load the ovens differently so that more loaves of bread can be baked at one time. This new process will require that the ovens be loaded by​ hand, requiring additional manpower. This is the only thing to be changed.  The bakery currently makes 1 comma 500 loaves per month. The pay will be ​$8 per hour for employees and each employee works 160 hours per month. Charles Lackey can also improve the yield by purchasing a new blender. The new blender will mean an increase in his investment. This new blender will mean an increase in his costs of ​$150 per​ month, but he will achieve the same new output​ (an increase to 2 comma 025.00​) as the change in labor hours. ​a) Current productivity for 640 work hours​ = nothing ​loaves/dollar ​(round your response to three decimal​ places). If Charles chooses to increase the number of work hours to 864 in order to employ the new oven loading​ technique, then the productivity is​ = nothing ​loaves/dollar ​(round your response to three decimal​ places). ​b) If Charles instead chooses to purchase a new blender​ (while holding labor constant at 640 hours at ​$8 per ​hour)​, then the productivity is​ = nothing ​loaves/dollar ​(round your response to three decimal​ places). ​c) By adding​ manpower, the percentage increase in productivity is nothing​% ​(enter your response as a percentage rounded to two decimal places and include a minus sign if​ necessary). By purchasing a new blender​ (while holding labor constant at 640 hours at ​$8 per ​hour)​, the percentage increase in productivity is nothing​% ​(enter your response as a percentage rounded to two decimal places and include a minus sign if​ necessary).
Answer #1

a)Output = 1500 loaves per month

There are 640 work hours per month and employees are Paid $8 per hour. So labor cost per month= 640 hours x $8 = $5120

Current productivity for 640 work hours = output / labor cost = 1500 loaves / $5120 = 0.293 loaves per dollar

The earlier Output was 1500 loaves per month which is increased by 35%. So now the output = 1500 + (35% of 1500) = 1500+525 = 2025 loaves

There are 864 work hours per month and employees are Paid $8 per hour. So labor cost per month= 864 hours x $8 = $6912

Prooductivity = output / labor cost = 2025 loaves / $6912 = 0.293 loaves per dollar

b) with the purchase of the new blender, the output is the same new output as the change in labor hours

The earlier Output was 1500 loaves per month which is increased by 35%. So now the output = 1500 + (35% of 1500) = 1500+525 = 2025 loaves

So output = 2025 loaves per month  

The per hour labor cost is $8.so for 640 labor hours, labor cost = $8 x 640 = $5120

The new blender increases the cost by $150 per month.

So now the total cost = labor cost + blender cost = $5120+$150 = $5270

So productivity = output / Total cost

= 2025 loaves / $5270

= 0.384 loaves per dollar

C) Before adding the manpower productivity = 0.293 oaves per dollar (calculated in part a)

After adding the manpower productivity = 0.293 oaves per dollar (calculated in part a)

% increase in productivity = (productivity after adding the manpower-productivity before adding the manpower) / productivity before adding the manpower

= (0.293-0.293) / 0.293

= 0/0.293

= 0%

Before purchasing the new blender, productivity = 0.293 loaves per dollar (calculated in part a)

After purchasing the new blender, productivity = 0.385 loaves per dollar (calculated in part b)

% increase in productivity = (productivity after purchasing the new blender-productivity before purchasing the new blender) / productivity before purchasing the new blender

= (0.384-0.293) / 0.293

= 0.091/0.293

= 0.3106

= 31.06%